You're about to buy a new car for $10,000. The dealer offers you a one-year loan where you pay $882.88 every month for the next 12 months. Since you pay $882.88 * 12 = $10,595 in total, the dealer claims that the loan's annual interest rate is (10,595-10,000)/10,000 = 5.95%.

What is the actual effective annual rate?

What rate should the dealer quote by law? (in decimals)
Answer #1

Annual effective rate (EAR) = (1+ APR/12)^12 -1

= (1+ 5.95%/12)^12 -1

= 6.114 %

PV of loan = $10,000

Payment every month (PMT) = $882.88

Number of months(nper) = 12

Rate needed to be quoted by dealer as per law (rate) = =RATE(12,-882.88,10000) RATE(nper, pmt, pv, [fv], [type] (guess]) = 0.9% per month = 10.8% pa